# ALPINE: Unveiling the Planning Capability of Autoregressive Learning in Language Models

### v2

@@ 3 Path-Finding Transformers: Expressive Power and Learning Capacity
For simplicity, we omit all layer normalizations in this construction.
Before presenting the detailed proof, we provide a summary of our construction. In essence, we utilize the attention layer to attend the output solely to the target node $t$. This approach allows the distribution of next token $u_{k+1}$ to become a function of both the current node $u_{k}$ and the target node $t$ (as formulated in Section 2). Then, by integrating the adjacency matrix $\bm{A}^{\text{\rm true}}$ into the MLP layer and the reachability matrix $\bm{R}^{\text{\rm true}}$ into the matrix ${\bm{W}}^{V}$ in the attention layer, we extract row vectors ${\bm{R}}^{\text{\rm true}}_{(t,:)}$ and ${\bm{A}}^{\text{\rm true}}_{(u_{k},:)}$ from ${\bm{R}}^{\text{\rm true}}$ and ${\bm{A}}^{\text{\rm true}}$, respectively, corresponding to the target node $t$ and current node $u_{k}$. Specifically, ${\bm{R}}^{\text{\rm true}}_{(t,:)}$ and ${\bm{A}}^{\text{\rm true}}_{(u_{k},:)}$ are stored by $\text{MHA}({\bm{H}}_{0})$ and $\text{FFN}(\text{MHA}({\bm{H}}_{0})+{\bm{H}}_{0})$, respectively. By selecting proper coefficients, we can ignore the effect of the remaining term ${\bm{H}}_{0}$ in $\text{Transformer}({\bm{H}}_{0})$ and only keep a weighted sum of ${\bm{R}}^{\text{\rm true}}_{(t,:)}$ and ${\bm{A}}^{\text{\rm true}}_{(u_{k},:)}$. Following the softmax layer, the non-negligible entries in the final vector correspond to the feasible next nodes. With this encoding, the Transformer serves as a simulator of the idealized Algorithm 1 with input $\bm{A}=\bm{A}^{\text{\rm true}}$ and $\bm{R}=\bm{R}^{\text{\rm true}}$.
We now provide the detailed proof. Suppose the input token sequence is “$s\ t\ s\ u_{1}\ u_{2}\ \ldots\ u_{k}$” with $k\geq 0$, where $s$ ($=u_{0}$) and $t$ are the tokens of the source and target nodes, respectively, and nodes $s,u_{1},\cdots,u_{k}$ form a path that can reach node $t$ in graph $\mathcal{G}$. Our objective is to construct a $1$-layer, $1$-head Transformer model that consistently generate an out-neighbor $u_{k+1}$ of $u_{k}$, enabling a path from $u_{k}$ to $t$ in $\mathcal{G}$.
@@
Then we construct the attention layer of our Transformer. We only have one head and let ${\bm{W}}^{K}=\left(\begin{matrix}{\bm{0}}_{(|\mathcal{V}|+2)\times(|\mathcal{V}|+1)}\mid{\bm{1}}_{(|\mathcal{V}|+2)\times 1}\end{matrix}\right)^{\top}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}^{Q}=\sqrt{d}\cdot{\bm{I}}_{d\times d}$. Then we can compute ${\bm{H}}_{0}{\bm{W}}^{K}=\left(\begin{matrix}{\bm{0}}_{(|\mathcal{V}|+2)\times 1}\mid c_{0}\cdot{\bm{1}}_{(|\mathcal{V}|+2)\times 1}\mid{\bm{0}}_{(|\mathcal{V}|+2)\times(k+1)}\end{matrix}\right)^{\top}$ (i.e. second rows are all $c_{0}$’s and other rows are all 00’s) and ${\bm{H}}_{0}{\bm{W}}^{Q}=\sqrt{d}\cdot{\bm{H}}_{0}$.
@@
where $c_{2}$ are positive parameters to be decided, and ${\bm{b}}_{2}={\bm{0}}$. By this way, we have
@@
Then we fix $c_{1}=c_{2}$ and let them be large enough. In this case, the dominant entries in $({\bm{H}}_{1})_{(k+3,:)}$ represent the nodes that are both the out-neighbor of $u_{j}$ and reachable to $t$, since those entries will have the value of $2c_{1}$ while others entries are at most $c_{1}+1$. This means that $({\bm{H}}_{1})_{(k+3,:)}$ can correctly indicates the next node $u_{k+1}$. Specifically, let ${\bm{W}}_{o}=\left(\begin{matrix}{\bm{I}}_{(|\mathcal{V}|+1)\times(|\mathcal{V}|+1)}\mid{\bm{0}}_{(|\mathcal{V}|+1)\times 1}\end{matrix}\right)^{\top}\in\mathbb{R}^{d\times M}.$ Then the final output approaches the following vector
$\lim_{c_{0},c_{1}=c_{2}\to\infty}\hat{u}_{k+1}$ $=\lim_{c_{0},c_{1}=c_{2}\to\infty}\text{softmax}(({\bm{H}}_{1})_{(k+3,:)}{\bm{W}}_{o})$
@@
where $C$ is the number of nodes that are both the out-neighbor of $u_{k}$ and reachable to $t$. Thus, this encoding guarantees that this is exactly the correct output of the next node. Hence, for any $\varepsilon>0$, we can always find a $1$-layer, $1$-head, and $(|\mathcal{V}|+2)$-embedding-size Transformer that provides the correct response with probability at least $1-\varepsilon$ by selecting large enough parameters $c_{0},c_{1},c_{2}$.
@@
Having established the mathematical existence of a Transformer model capable of accomplishing path finding in a given network, as demonstrated in Theorem 2, we now shift our focus to the following fundamental question
Can the Transformer architecture, trained on sufficient path data with an auto-regressive loss as in Equation (7) and using the gradient descent (GD) method, learn the adjacency and reachability matrices and carry out path finding similar to the idealized Algorithm 1?
Through a combination of theoretical analysis and empirical evaluation presented in the following section, our primary investigation aims to address the aforementioned question.
First, it is important to note that the Transformer may not be capable to learn the exact true adjacency and reachability matrices of the underlying graph. Instead, it can only learn the relevant information that is directly encoded in the observed training data $\mathcal{D}$. Therefore, we define the observed adjacency and reachability matrices based on the training data $\mathcal{D}$ as follows.
$\bm{A}^{\text{obs}}_{(i,k)}(\mathcal{D})=\begin{cases}1,&\text{if}\exists\bm{u}\in\mathcal{D},n\in[3,N-1]\text{ s.t. }u_{n}=i,u_{n+1}=k\\ 0,&\text{otherwise}\end{cases}$
$\bm{R}^{\text{obs}}_{(t,k)}(\mathcal{D})=\begin{cases}1,&\text{if}\exists\bm{u}\in\mathcal{D},n\in[4,N]\text{ s.t. }u_{2}=t,u_{n}=k\\ 0,&\text{otherwise}.\end{cases}$
Naturally, the observed adjacency matrix $\bm{A}^{\text{obs}}(\mathcal{D})$ only records the edges $(i,k)$ that appears in some path within the training data $\mathcal{D}$. On the other hand, the observed reachability matrix $\bm{R}^{\text{obs}}(\mathcal{D})$ exhibits more nuanced distinctions from the true reachability matrix. It only records that $t$ is reachable from node $k$, if the training data $\mathcal{D}$ contains a path (sequence) whose destination is $t$ and $k$ appears as a non-source node on the path. We call such pairs $(t,k)$ observed reachable pairs. Therefore, the observed reachability matrix would miss the following types of reachable pairs $(t,k)$ in $\mathcal{G}$ (referred as non-observed reachable pairs): (i) there is no path in $\mathcal{D}$ that contains a sub-path from $k$ to $t$, even if a path from $k$ to $t$ can be obtained by concatenating several sub-paths appeared in $\mathcal{D}$; (ii) there are some paths in $\mathcal{D}$ that contains a sub-path from $k$ to $t$, however, $t$ is not the target node in these paths; (iii) there are some paths in $\mathcal{D}$ that contains a sub-path from $k$ to $t$ and $t$ is the target node in these paths, however, $k$ is always the source node of these paths.
In Section 4, we show that in a simplified Transformer model, the learning is limited to the observed adjacency and reachability matrices, rather than the true underlying matrices. The following presents an informal version of the result:
Theorem 3 (Learning Capacity and Limitation of Path-Finding Transformer: An Informal version).
By using auto-regressive loss and training with gradient descent, a simplified Transformer architecture with $1$-layer, $1$-head, and $O(|\mathcal{V}|)$-embedding-size simulates Algorithm 1 with $\bm{R}=\bm{R}^{\text{obs}}(\mathcal{D})$ and $\bm{A}=\bm{A}^{\text{obs}}(\mathcal{D})$.
The formal analytical result is presented as Theorem 4, which gives captures the direction of changes of the parameters in the learnable matrices of the simplified Transformer when following the gradient descent calculation. We then discuss that how this result indicates that the simplified Transformer effective learns the observed adjacency matrix and the observed reachability matrix, and its inference procedure indeed align with the workings of Algorithm 1. Specifically, in the simplified Transformer, the observed adjacency matrix is encoded within the weights of the feed-forward network (FFN) as illustrated in Figure 1, while the observed reachability matrix is encoded in the value matrix, as depicted in Figure 2.
Next, in Section 5, we present the results of our empirical evaluation which is based on extensive experiments. We report the accuracy achieved by the Transformer models with various hyperparameters (Figure 3). Furthermore, we provide visualizations that demonstrate the Transformer’s ability to learn attention (Figure 4) as well as the information about adjacency and reachability matrices (Figures 5 and 6). Notably, our findings reveal that even a large Transformer model fails to learn the reachability matrix beyond $\bm{R}^{\text{obs}}(\mathcal{D})$, resulting poor path-finding accuracy for those unobserved reachable pairs (Figure 6). To further validate our approach, we conduct experiments on a realistic planning dataset called Blockswords. The accuracy, attention, adjacency matrix and reachability matrix are shown in Figure 7 and Figure 8. Importantly, our empirical results align closely with our theoretical findings discussed in Section 4.
4 Gradient-based Analysis for Path Finding
Let $\mathcal{D}$ be the path dataset as described in Section 2.2. In this section, we show analytically that even with only one layer and one head, the Transformer architecture could learn both the adjacency matrix and the reachability matrix from the dataset $\mathcal{D}$ and then predict the next node on a path, similar to what is done in Algorithm 1. Let $N_{i,j,k}$ be the number of times in $\mathcal{D}$ that i) the current node is $i$; ii) the destination node is $j$ and iii) the next node is $k$, and let $N_{i,j}=\sum_{k}N_{i,j,k}$.
[…]
  • The attention weight is only on the target node (the second token), i.e., we manually set every row in $\textbf{softmax}\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d_{k}}}\right)$ (in Eq. (1)) to be a one-hot vector with the second coordinate being $1$. Moreover, we set the positional embedding matrix $\bm{W}_{p}=\bm{0}$, since it is usually used to adjust the attention weights.
  • We remove all the non-linear layers (e.g., the layer normalizations), and use $\text{FFN}(\bm{X})=\bm{X}\bm{W}^{M}$ instead of Eq. (3), and use $\text{Transformer}(\bm{X})=\text{FFN}(\bm{X})+\text{MHA}(\bm{X})$ instead of Eq. (4).
  • The token embedding matrix $\bm{W}_{t}$ and the output weight matrix $\bm{W}_{o}$ are set to be identity, i.e., $\bm{W}_{t}=\bm{W}_{o}=\bm{I}$.
Since there is only one layer and one head, for simplicity, we use $\bm{W}^{V}$ to represent the weight of the value matrix in the attention layer. Under the above Transformer structure,
$(\bm{H}_{L})_{(n,:)}\bm{W}_{o}=(\bm{U}\bm{W}_{t}\bm{W}^{M}+\bm{\alpha}\bm{U}\bm{W}_{t}\bm{W}^{V})_{(n,:)}\bm{W}_{o}=(\bm{U}\bm{W}^{M}+\bm{\alpha}\bm{U}\bm{W}^{V})_{(n,:)}=\bm{W}^{M}_{(u_{n},:)}+\bm{W}^{V}_{(u_{2},:)},$ (8)
where $\bm{\alpha}$ is the manually set attention weight matrix (every row is a one-hot vector with the second coordinate being $1$). Therefore, the weight vector when predicting the $(n+1)^{th}$ token is $\textbf{softmax}(\bm{W}^{M}_{(u_{n},:)}+\bm{W}^{V}_{(u_{2},:)})$, and the prediction probability is
$\hat{u}_{n+1,k}=\frac{\exp(\bm{W}^{M}_{(u_{n},k)}+\bm{W}^{V}_{(u_{2},k)})}{\sum_{\ell}\exp(\bm{W}^{M}_{(u_{n},\ell)}+\bm{W}^{V}_{(u_{2},\ell)})}.$ (9)
We prove the following theorem.
Theorem 4.
Under the cross-entropy loss $\ell(\mathcal{D})$, for all possible $(i,k)$ pairs, i) if $\sum_{j}N_{i,j}=0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is always 0; ii) if $\sum_{j}N_{i,j}>0$ but $\sum_{j}N_{i,j,k}=0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is always positive; iii) if $\sum_{j}N_{i,j,k}>0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is negative when $\bm{W}^{M}_{(i,k)}$ converges to $\infty$- ∞. Similarly, for all possible $(j,k)$ pairs, i) if $\sum_{i}N_{i,j}=0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{V}_{(j,k)}}$ is always 0; ii) if $\sum_{i}N_{i,j}>0$ but $\sum_{i}N_{i,j,k}=0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{V}_{(j,k)}}$ is always positive; iii) if $\sum_{i}N_{i,j,k}>0$, then ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{V}_{(j,k)}}$ is negative when $\bm{W}^{V}_{(j,k)}$ converges to $\infty$- ∞.
@@ 4 Gradient-based Analysis for Path Finding
We only prove the first part of this theorem, since the proof of the second part is almost the identical.
@@
In case i), $\sum_{j}N_{i,j}=0$ implies that $\sum_{j}N_{i,j,k}=0$. Hence ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is always zero.

### v3

@@ Appendix B Proof of Theorem 2
For simplicity, we omit all layer normalizations in this construction. Suppose the input token sequence is “$s\ t\ s\ u_{1}\ u_{2}\ \ldots\ u_{k}$” with $k\geq 0$, where $s$ ($=u_{0}$) and $t$ are the tokens of the source and target nodes, respectively, and nodes $s,u_{1},\cdots,u_{k}$ form a path that can reach node $t$ in graph $\mathcal{G}$. Our objective is to construct a $1$-layer and $1$-head Transformer model that generates an out-neighbor $u_{k+1}$ of $u_{k}$ such that there exists at least one path from $u_{k+1}$ to $t$ in $\mathcal{G}$.
In essence, we utilize the attention layer to attend the output solely to the target node $t$. This approach allows the distribution of next token $u_{k+1}$ to become a function of both the current node $u_{k}$ and the target node $t$ (as formulated in Section 2). Then, by integrating the adjacency matrix $\bm{A}^{\text{\rm true}}$ into the FFN layer and the reachability matrix $\bm{R}^{\text{\rm true}}$ into the matrix ${\bm{W}}^{V}$ in the attention layer, we extract row vectors ${\bm{R}}^{\text{\rm true}}_{(t,:)}$ and ${\bm{A}}^{\text{\rm true}}_{(u_{k},:)}$ from ${\bm{R}}^{\text{\rm true}}$ and ${\bm{A}}^{\text{\rm true}}$, respectively, corresponding to the target node $t$ and current node $u_{k}$. By selecting proper coefficients, we can let the output be the sum of ${\bm{R}}^{\text{\rm true}}_{(t,:)}$ and ${\bm{A}}^{\text{\rm true}}_{(u_{k},:)}$. Following the softmax layer, the non-negligible entries in the final vector correspond to the feasible next nodes. With this encoding, the Transformer serves as a simulator of Algorithm 1 with input $\bm{A}=\bm{A}^{\text{\rm true}}$ and $\bm{R}=\bm{R}^{\text{\rm true}}$.
@@
Then we construct the attention layer of the Transformer. We only have one head and let ${\bm{W}}^{K}=\left(\begin{matrix}{\bm{0}}_{(|\mathcal{V}|+2)\times(|\mathcal{V}|+1)}\mid{\bm{1}}_{(|\mathcal{V}|+2)\times 1}\end{matrix}\right)^{\top}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}^{Q}=\sqrt{d}\cdot{\bm{I}}_{d\times d}$. Then we can compute ${\bm{H}}_{0}{\bm{W}}^{K}=\left(\begin{matrix}{\bm{0}}_{(|\mathcal{V}|+2)\times 1}\mid c_{0}\cdot{\bm{1}}_{(|\mathcal{V}|+2)\times 1}\mid{\bm{0}}_{(|\mathcal{V}|+2)\times(k+1)}\end{matrix}\right)^{\top}$, i.e., second rows are all $c_{0}$’s and other rows are all 00’s, and ${\bm{H}}_{0}{\bm{W}}^{Q}=\sqrt{d}\cdot{\bm{H}}_{0}$.
@@
where $c_{2}$ is a positive parameter to be decided, and ${\bm{b}}_{2}={\bm{0}}$. By this way, we have
@@
Then we fix $c_{1}=c_{2}$ and let them be large enough. In this case, the dominant entries in $({\bm{H}}_{1})_{(k+3,:)}$ represent the nodes that are both the out-neighbor of $u_{j}$ and reachable to $t$, since those entries will have the value of $2c_{1}$ while other entries are at most $c_{1}+1$. This means that $({\bm{H}}_{1})_{(k+3,:)}$ can correctly indicates the next node $u_{k+1}$. Specifically, let ${\bm{W}}_{o}=\left(\begin{matrix}{\bm{I}}_{(|\mathcal{V}|+1)\times(|\mathcal{V}|+1)}\mid{\bm{0}}_{(|\mathcal{V}|+1)\times 1}\end{matrix}\right)^{\top}\in\mathbb{R}^{d\times M}.$ Then the final output approaches the following vector
$\lim_{c_{0},c_{1}=c_{2}\to\infty}\hat{\bm{u}}_{k+1}$ $=\lim_{c_{0},c_{1}=c_{2}\to\infty}\text{softmax}(({\bm{H}}_{1})_{(k+3,:)}{\bm{W}}_{o})$
@@
where $C$ is the number of nodes that are both the out-neighbor of $u_{k}$ and reachable to $t$. Thus, this encoding guarantees that for any $\varepsilon>0$ and $Q>0$, we can always construct a $1$-layer, $1$-head, and $(|\mathcal{V}|+2)$-embedding-size Transformer that provides the correct next token with probability at least $1-\frac{\varepsilon}{2Q}$ by selecting large enough parameters $c_{0},c_{1},c_{2}$.
Then we prove that there exists a $Q$ such that this Transformer can output a correct path for every valid source and target node pair with probability at least $1-\varepsilon$. Suppose we can output all the nodes that are both the out-neighbor of the current node and reachable to the target node with the same probability in each round without any error. Then, whatever the current node is, there is a probability of at least $\frac{1}{|\mathcal{V}|^{|\mathcal{V}|}}$ that the target node is reached within the next $|\mathcal{V}|$ generated nodes. Therefore, the target node is reached in $c_{3}\cdot|\mathcal{V}|$ steps with probability at least $1-\left(1-\frac{1}{|\mathcal{V}|^{|\mathcal{V}|}}\right)^{c_{3}}$, where $c_{3}\in\mathbb{N}$ is a positive integer. We let $c_{3}=\log_{1-\frac{1}{|\mathcal{V}|^{|\mathcal{V}|}}}\frac{\epsilon}{2}$ and $Q=c_{3}\cdot|\mathcal{V}|$. Then, according to the Union bound, the Transformer can output a correct path in $Q$ steps with an error rate of at most $\frac{\varepsilon}{2Q}\cdot Q+\frac{\varepsilon}{2}=\varepsilon$.
@@
Appendix C Proof of Theorem 3
See 3
@@ Appendix C Proof of Theorem 3
We only prove the first part of this theorem, since the proof of the second part is almost identical.
@@
In case i), $\sum_{j}N_{i,j}=0$ implies that $\sum_{j}N_{i,j,k}=0$. Hence ${\partial\ell(\mathcal{D})\over\partial\bm{W}^{M}_{(i,k)}}$ is always 0.